# Differentially Private Clustering in Data Streams

### v2

@@ 1.2 Our Results
    (Accuracy) With high probability,
    $\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})$ $\leq O_{\gamma}(1)\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})+O_{\gamma}(1)\cdot V^{\prime}(d,k,\varepsilon,\delta,T,\Lambda)$ (2)
    where ${\mathcal{S}}$ denotes the set of all points given by the stream before timestamp $t$, ${\mathcal{C}}_{{\mathcal{S}}}^{opt}$ is the optimal solution, and $V^{\prime}(d,k,\varepsilon,\delta,T,\Lambda)=\tilde{O}(\Lambda^{2}\cdot\operatorname*{poly}(\log(T),\log(1/\delta),d,\frac{1}{\varepsilon},k))$.
@@
    (Space) ${\mathcal{A}}^{\prime}$ consumes $\tilde{O}(k^{1.5}\cdot\operatorname*{poly}(\log(T),\log(\frac{1}{\delta}),d,\frac{1}{\varepsilon}))$.
The notation $O_{x}(\cdot)$ ignores factors involving $x$.
Remark 1.
The space complexity achieved by our framework depends on the additive error of the offline DP semicoreset (see Definition 3 for the definition of semicoreset) algorithm used as a black-box. In the case of Theorem 2, the DP $k$-means algorithm from [42] achieves an additive error with $k^{1.5}$ dependency whereas their DP coreset algorithm has $\operatorname*{poly}(k)$ dependency. Thus we use their DP $k$-means algorithm to compute a DP semicoreset albeit with a slight loss in privacy to save on space (see Theorem 7 for the transformation). In the case of Theorem 1, the additive error of the DP $k$-means clustering algorithm and DP coreset from [29] has the same dependency in $k$. Hence we use their DP coreset algorithm directly.
Finally, note that if there exists a DP semicoreset algorithm which achieves a better additive error in the offline setting, then our framework immediately gives a DP clustering algorithm with better space bounds in the continual release setting.
1.3 Comparison with Concurrent Work [35]
[35] concurrently released results on differentially private $k$-means clustering under continual observation. However their work is mostly incomparable with ours: their main focus is to optimize the approximation ratio under the setting of continual observation, but their algorithm uses a large space. In contrast, our main focus is to optimize the approximation ratio using as small space as possible, and the property that our algorithm is DP under continual observation (or equivalently, continual release) setting is a side outcome of our algorithm.
1.4 Technical Overview
Our techniques apply to both $k$-means and $k$-median clustering, but we assume we are dealing with $k$-means for simplicity. Before discussing our algorithm in more detail, we first outline the challenges to designing a DP $k$-means clustering algorithm in the continual release setting.
Vanilla Merge and Reduce approaches fail.
A natural idea is to follow a standard approach called Merge-and-Reduce [27] in the streaming clustering literature. A coreset of a set of points is a weighted subset of the points such that the clustering cost of the coreset with respect to any $k$-centers $C$ is a good approximation of the clustering cost of the original point set with respect to $C$. We refer readers to Definition 2 for a formal definition of coreset. According to [18], any point set has a $(1+\gamma)$-coreset with size $k\cdot\operatorname*{poly}(1/\gamma)$. Merge-and-Reduce approach suggests the following. We maintain $C_{1},C_{2},\cdots,C_{L}$ for $L=\log(T)$ during the stream such that $C_{1}\cup C_{2}\cup\cdots\cup C_{L}$ is a $(1+\gamma)$-coreset of the points already arrived. We choose some threshold $M=k\operatorname*{poly}(\log(T))$. When a new point $x$ comes, we make $C_{1}\leftarrow C_{1}\cup\{x\}$. Then we do the following check: if there is $i\in[L]$ such that $|C_{i}|\geq M$, we compute a $(1+1/\operatorname*{poly}(\log(T)))$-coreset $C^{\prime}$ of $C_{i}$, and we set $C_{i}\leftarrow\emptyset$ and $C_{i+1}\leftarrow C_{i+1}\cup C^{\prime}$. We repeat the above checking process until every $C_{i}$ has size $|C_{i}|\leq M$. It is clear that at any time, $C_{i}$ is a coreset of a subset $S_{i}$ of input points and $S_{1}\cup S_{2}\cup\cdots\cup S_{L}$ is the entire set of the input points. By induction, we can show that $C_{i}$ must be a $(1+1/\operatorname*{poly}(\log(T)))^{i}$-coreset of $S_{i}$, and thus $C_{1}\cup C_{2}\cup\cdots\cup C_{L}$ is a $(1+1/\operatorname*{poly}(\log(T)))^{L}=(1+O(\log(T))/\operatorname*{poly}(\log(T)))$-coreset of the entire input point set.
However, the above approach fails when using DP coreset for $C_{1}$. The reason is that any DP coreset must introduce some additive error $\eta=\Omega(\Lambda^{z})$ (see e.g., [29]). In particular, let us follow the above Merge and Reduce process, but when $C_{1}\geq M$, we compute a DP coreset $C^{\prime}$ of $C_{1}$ instead. Then $C^{\prime}$ is merged to $C_{2}$, and thus $C_{2}$ is a coreset of $S_{2}$ with some additive error $\eta$. Now consider $\hat{C}$ is a corest of $\hat{S}$ with additive error $\hat{\eta}$ and $\bar{C}$ is a corest of $\bar{S}$ with additive error $\bar{\eta}$. Then even if we compute any non-DP coreset $\tilde{C}$ of $\hat{C}\cup\bar{C}$ with no relative error, $\tilde{C}$ is a coreset of $\hat{S}\cup\bar{S}$ with additive error $\bar{\eta}+\hat{\eta}$. The above example implies that the additive error can actually accumulate. Therefore, $C_{i}$ is a coreset of $S_{i}$ with additive error $2^{O(i)}\cdot\Omega(\Lambda^{z})$ which can be as large as $T^{\Omega(1)}\cdot\Lambda^{z}$.
To overcome the above issue, we develop a novel method of DP Merge and Reduce: instead of running Merge and Reduce for all points together, we firstly partition the space $\mathbb{R}^{d}$ such that close points are likely to be in the same group, then we run Merge and Reduce in a DP way for points in each group. In this way, we are able to show that the additive error introduced can always be bounded by the optimal clustering cost with some small multiplicative factors. In the remainder of this section, we briefly discuss the high level ideas of our approach.
Our Approach.
For every timestamp $t\in[T]$, our algorithm can be split into the following main steps — (1) Compute a set of centers ${\mathcal{F}}$ in an online DP fashion that satisfies a bicriteria approximation111a bicriteria approximation algorithm in this case means that the algorithm may output a set of more than $k$ centers to $k$-means. (2) Maintain DP semicoresets of the points assigned to centers in ${\mathcal{F}}$ in parallel via multiple DP Merge-and-Reduce instances (3) Output the union of these semicoresets called $\hat{{\mathcal{Y}}}$. In a post-processing step — Compute a non-DP $k$-means $\rho$-approximation algorithm on $\hat{{\mathcal{Y}}}$.
@@ 1.4 Technical Overview
The bicriteria approximation uses two main ingredients — quadtrees and heavy hitters. A quadtree creates a nested series of grids that partitions ${\mathbb{R}}^{d}$ and can be used to embed input points into a Hierarchically Separated Tree (HST) metric, which often makes computing $k$-means cost simpler. We use this embedding to map every input point to the center of a grid (or cell) at every quadtree level. For a fixed level, our goal is to approximately choose the $O(k)$ cells that have the most points, i.e., we want to find the “heaviest” cells in a DP fashion and store them as candidate centers in set ${\mathcal{F}}$. We achieve this by hashing the cells into $O(k)$ substreams and running a continual release black-box DP heavy hitter algorithm from [24] on each hash substream. Since with large enough probability, the heavy cells will not collide, this achieves our goal. Note that since we need to do this over logarithmically many levels of the quadtree, we will end up with a bicriteria approximation.
Theorem 3 (Bicriteria approximation).
There exists an $\varepsilon$-DP algorithm that for every timestamp $t\in[T]$, computes a weighted set of $O({k}\log^{2}(k)\log(\Lambda)\log T)$ centers with $d^{O(1)}$-multiplicative error to the best $k$-means (or $k$-median) clustering and $\tilde{O}(\frac{k\rho\Lambda^{2}}{\varepsilon}\cdot(d\log T)^{O(1)}))$-additive error in $O(k\log^{2}\Lambda\log(k)\operatorname*{poly}\left(\log\left({T}{\Lambda}k\right)\right))$ space.
Reducing Multiplicative Error.
At this point, one idea is to assign input points to candidate centers in ${\mathcal{F}}$ obtained from the bicriteria approximation in an online fashion and release a weighted DP semicoreset at every timestep. The postprocessing step of applying an offline non-DP clustering algorithm on this weighted DP coreset would ensure a DP set of $k$ centers, however the cost of clustering of the resulting set of centers incur a multiplicative error of $d^{O(1)}$ (see Theorem 3), which can be quite large especially in high dimensional Euclidean space.
@@
Inspired by [17] who recently used techniques introduced by [14] to obtain a near optimal DP Clustering algorithm in the MPC setting, we define rings $R_{r}$ ( where $1\leq r\leq\log\Lambda$) based on distance from the current centers in set ${\mathcal{F}}$ and map input points to these rings. The main idea is to compute DP semicoresets for points in each ring and then by taking the union of the semicoresets over all rings, we can obtain a DP semicoreset for the stream of points seen so far. Since the DP semicoreset algorithms we use [29, 42] achieve constant multiplicative error, intuitively, this technique will result in constant multiplicative error of the resulting DP semicoreset. In particular, to compute DP semicoresets for points in each ring we run Merge and Reduce in a DP manner.
DP Merge and Reduce.
We design an online differentially private variant of the celebrated Merge and Reduce framework [33, 1] to compute DP semicoresets for points in each ring $R_{r}$ ( where $1\leq r\leq\log\Lambda$). An informal overview of the (non-DP) Merge and Reduce technique is — partition the input stream into blocks, compute a coreset for each block, take the union of the resulting coresets (merge step), and compute the coreset of the union (reduce step) in a tree-like fashion. On a high-level, our framework computes DP semicoresets at the base level (of the tree) using a DP semicoreset algorithm (e.g. [42, 29]) and then computes coresets for subsequent levels using a non-DP coreset Algorithm (e.g. [19]). Thus using our DP Merge and Reduce framework, at any given timestep, the number of input points we actually need to store is roughly equal to the size of the block at the base level or the threshold (denoted as $M$). This is because as soon as the number of input points exceed $M$, we can compute a DP coreset for this block. We perform this check privately via the sparse vector technique [23]. Note that the size of the base level block controls the height of the tree, i.e., larger $M$ results in smaller height of the tree which in turn results in smaller additive error of the resulting DP coreset. However, we show that we can store a sublinear number of input points and achieve reasonable additive error for a fixed instance of DP Merge and Reduce (see Theorem 10 for a formal statement).
@@
Although we show that a fixed instance of DP Merge and Reduce for a fixed ring releases a DP semicoreset that achieves reasonably small additive error $\approx\tilde{O}(\frac{N_{r}}{M}+\log(T))$ where $N_{r}$ is the number of points in ring $R_{r}$, the additive error of the union of these DP semicoresets over all rings can be large, e.g., linear in $T$. We use ideas similar to [17] to address this challenge — by using the properties of the bicriteria solution and carefully choosing threshold $M$ for the DP Merge and Reduce instances, we show that we can charge the large additive error obtained by summing up the cost of clustering over all rings to a small constant-factor blow up in the multiplicative error (see Theorem 12). In particular, we achieve an additive error that is $\approx\tilde{O}(\operatorname*{poly}(\log T))$.
@@
Finally, observe that the bicriteria solution ${\mathcal{F}}$ can change over time, i.e., new centers can be added to ${\mathcal{F}}$. Thus whenever a new center (or centers) are added to ${\mathcal{F}}$, we redefine the rings (based on distance to the set ${\mathcal{F}}$) and initiate new DP Merge and Reduce instances for each ring. See Algorithm 1 for more details.
2 Preliminaries

### v3

@@ 1.4 Our Techniques
    Compute a set of centers $\mathcal{F}$} in an online DP fashion that satisfies a bicriteria approximation (algorithm 3) to $k$-means.
@@
    Maintain DP semicoresets of the points assigned to centers in $\mathcal{F}$} in parallel via multiple DP Merge-and-Reduce (algorithm 4) instances
@@
    Output the union of these semicoresets called $\hat{{\mathcal{Y}}}$.
In a post-processing step — Compute a non-DP $k$-means $(1+\gamma)$-approximation algorithm on the output $\hat{{\mathcal{Y}}}$.
In the remainder of this section, we briefly discuss the different components of our approach.
@@
We first design a sublinear space online $\varepsilon$-DP algorithm that outputs a set of $\tilde{O}(k\log(T))$ candidate centers satisfying a $d^{O(1)}$-multiplicative approximation and $\tilde{O}(\frac{d^{2}\Lambda^{2}k}{\varepsilon}\cdot\operatorname*{poly}\log(T))$ additive error. The bicriteria approximation algorithm uses two main ingredients — quadtrees and heavy hitters. A quadtree creates a nested series of grids that partition $\mathbb{R}^{d}$ and can be used to embed input points into a Hierarchically Separated Tree (HST) metric, which often simplifies the analysis of $k$-means cost. We initialize our quadtree with $\log(\Lambda)$ levels, and use this embedding to map every input point to the center of a grid cell at each level.
For a fixed level, our goal is to approximately choose $\tilde{O}(k)$ cells that have the most points and store them as candidate centers in set $\mathcal{F}$} while using low space. To do this, we first hash cells of the fixed quadtree level to $w=O(k)$ buckets, and at every timestep, we track the DP size of each bucket through the standard Binary Mechanism [24, 9], as well as privately compute the heavy hitters of each bucket using the DP heavy hitter algorithm from [27] . Note that we need to track the DP size of each bucket using the Binary Mechanism separately to ensure that we can prune false positives among candidate centers. Without this pruning step, cells with very small counts could be mistakenly retained as heavy hitters. To the best of our knowledge this is the first use of DP Heavy hitters for DP clustering in the continual release setting.
Finally, the algorithm outputs the cumulative set of candidate centers $\mathcal{F}$} across each of the $\log\Lambda$ levels of the quadtree, yielding a bicriteria approximation. See Theorem 4 for a formal statement.
Grouping points and applying Merge and Reduce.
Instead of running a single instance of the DP Merge and Reduce framework over the entire stream (as described in the naive approach), our strategy is to partition the space ${\mathbb{R}}^{d}$ into groups such that input points close to a candidate center (of the bicriteria solution $\mathcal{F}$}) are in the same group and then run $\log(\Lambda)$ parallel instances of the DP Merge and Reduce framework per group. We use a partitioning technique introduced by [15] and define rings $R_{r}$ (where $1\leq r\leq\log\Lambda$) based on their distance from the current centers in $\mathcal{F}$} and map input points to these rings (see Definition 3 for a formal definition).
The intuition is that any DP coreset necessarily incurs additive error that scales with the square of the diameter of the set it is applied to. If we ran a single DP Merge and Reduce on all points assigned to a center, the diameter could be as large as $\Lambda$, leading to $\Omega(\Lambda^{2})$ additive error. By splitting the group into $\log(\Lambda)$ concentric rings, each ring has diameter only $O(2^{r})$, so the additive error of the DP routine on ring $R_{r}$ is $O(2^{2r})$. This error is now proportional to the cost contributed by points in that ring, which allows us to keep the total error under control.
@@
Now we can run the DP Merge and Reduce framework and obtain DP semicoresets for points in each ring, and by taking the union of these semicoresets over all rings, we obtain a DP semicoreset for the stream of points seen so far. Since the DP clustering algorithms we use as a black-box in our Merge and Reduce framework from [32, 47] achieve constant multiplicative error, this technique will result in constant multiplicative error of the resulting DP semicoreset.
@@
Because we are taking the union of DP semicoresets over all rings, the additive error resulting from this union depends on $\sum_{r}N_{r}/M$, where $N_{r}$ is the number of points in ring $r$ and $M$ controls when the DP Merge and Reduce framework is applied. At first glance, this sum could be as large as $T/M$, which seems too large to be useful.
The crucial insight is that we use a structural property of the bicriteria solution [^2] Intuitively, this property holds because the bicriteria solution already “pays” for the squared distances that define the rings: every point is assigned to a nearby candidate center, and the radius of its ring is proportional to this distance. (Lemma 22): the total sum of squared ring radii over all points is bounded, up to constants, by the $k$-means cost of the bicriteria solution $F$ (i.e., $\text{cost}(F,S)=\sum_{x\in S}d(x,F)^{2}$), which in turn is within a constant factor of the optimal $k$-means cost plus a small additive term. This property ensures that the additive error accumulated by the DP Merge and Reduce instances is controlled by the clustering cost. In turn, this allows us to charge the additive error to the multiplicative error, so that with an appropriate choice of $M$ the overall error is just a small constant-factor blow-up in the approximation plus a polylogarithmic additive term (see Section D.2 for details).
@@
Finally, observe that the bicriteria solution $\mathcal{F}$} can change over time, i.e., new centers can be added to $\mathcal{F}$}. Thus whenever a new center (or centers) are added to $\mathcal{F}$}, we redefine the rings (based on distance to the set $\mathcal{F}$}) and initiate new DP Merge and Reduce instances for each ring.
Comparison of our techniques to [18].
[18] gave DP $k$-means and $k$-median algorithms in the Massively Parallel Computation (MPC) model which achieved $O(1)$-multiplicative and $\operatorname*{poly}(k,d)$ additive error. Although their techniques are similar to ours, in the sense that they reduce the problem of solving DP clustering on the entire dataset to solving multiple instances of the DP clustering problem on partitioned input data, there are several differences in our work and new challenges that we have to deal with as a result.
First, their work does not tackle sublinear space in the continual release setting. Second, our bicriteria approximation algorithm uses a black-box continual release DP heavy hitters algorithm from [27] to find the candidate centers and thus requires a completely different analysis compared to the greedy approach used in [18]. More importantly, we have to implement a novel DP Merge and Reduce framework that works with both an offline DP clustering or DP coreset algorithm, which was not required in the MPC setting.
2 Differentially Private Clustering Framework
In this section, we first present our differentially private clustering framework in the continual release setting which is given by algorithm 1 and discuss the different components of the algorithm as well as its analysis in detail (we defer some of the proofs to Appendix D). Notably, our framework allows us to plug in any existing offline DP coreset or clustering algorithm to obtain a corresponding DP clustering algorithm in the continual release setting. The proofs of Theorem 1, Theorem 2 and Theorem 3 are stated in Section 2.1.
Data: Stream $\mathcal{S}$} of points $x_{1},\ldots,x_{T}\in{\mathbb{R}}^{d}$, Privacy parameters $\varepsilon,\varepsilon_{1},\delta_{1}$
Call $\textbf{Initialize}(\varepsilon)$ of BicriteriaDPCenters;
$\triangleright$ See algorithm 2
Initialize bicriteria solution $\mathcal{F}$$\emptyset$;
Initialize flag for new centers being added $flag_{new}=0$, and DP semicoreset $\hat{{\mathcal{Y}}}$$\emptyset$;
for new point $x_{t}$ do
 ${\mathcal{F}}_{t}\leftarrowUpdate$$(x_{t})$ of BicriteriaDPCenters;
 if ${\mathcal{F}}_{t}\neq\emptyset$ and $|{\mathcal{F}}\cap{\mathcal{F}}_{t}|<|{\mathcal{F}}_{t}|$ then
    $\triangleright$ New centers added to $\mathcal{F}$} — need to redefine rings
    $flag_{new}=1$;
 ${\mathcal{F}}\leftarrow{\mathcal{F}}\cup{\mathcal{F}}_{t}$;
 for $1\leq r\leq\log\Lambda$, run the following in parallel do
     Let $R_{r}$ represent the ring centered at $\mathcal{F}$} (see Definition 3);
    if $flag_{new}=1$ then
         Create new instance $\textbf{DP-Merge-Reduce}_{R_{r}}$ by calling $\textbf{Initialize}(\varepsilon,\varepsilon_{1},\delta_{1})$ of DP-Merge-Reduce;
       $\triangleright$ See algorithm 4
    if $x_{t}\in R_{r}$ then
[…]
    else
       $\hat{{\mathcal{Y}}}_{r}\leftarrow$ Update()($\perp$) of $\textbf{DP-Merge-Reduce}_{R_{r}}$;
 $flag_{new}=0$;
  Set $\hat{{\mathcal{Y}}}\leftarrow\hat{{\mathcal{Y}}}\cup(\cup_{r}\hat{{\mathcal{Y}}}_{r})$;
  Run non-DP $(1+\gamma$)-coreset algorithm on $\hat{{\mathcal{Y}}}$ and store this semicoreset as the new $\hat{{\mathcal{Y}}}$;
  Output the semicoreset $\hat{{\mathcal{Y}}}$;
  Delete existing input points (if any) in memory;
Algorithm 1 Main Algorithm
Before presenting the main algorithm, we formally define a ring centered at a set as we will use this to partition the space ${\mathbb{R}}^{d}$ in our algorithm.
Definition 3 (Ring centered at a Set).
Main Algorithm (algorithm 1).
When a new point $x_{t}$ arrives, our algorithm does the following
  1. 1.
    Update the bicriteria solution $\mathcal{F}$} (see algorithm 2)
  2. 2.
    If new centers have been added to $\mathcal{F}$} then create rings according to Definition 3 for the set $\mathcal{F}$} and add $x_{t}$ to a ring [^3] Notice that in the pseudo code algorithm 1 the symbol $\perp$ represents an empty update that is effectively ignored. This is needed for technical reasons to ensure DP by avoiding the value of the input affecting the number of events in the sub-streams.. For each ring $1\leq r\leq\log(\Lambda)$ create an instance of $\textsf{DP-Merge-Reduce}_{r}$ (see algorithm 4) which outputs a DP semicoreset per ring.
    If no new centers have been added to $\mathcal{F}$} in this timestep, then instead of creating new rings, the algorithm adds $x_{t}$ to an existing ring (and corresponding DP-Merge-Reduce instance).
  3. 3.
    Release the union of these DP semicoresets as $\hat{{\mathcal{Y}}}$.
In order to keep our space usage small, we apply a $(1+\gamma)$-approximation non-DP coreset algorithm to the union of semicoresets of these rings as $\hat{{\mathcal{Y}}}$.
Analysis.
We first state the theoretical guarantees of BicriteriaDPCenters and DP-Merge-Reduce as we need these results to state the guarantees of our Main Algorithm (see algorithm 1). The proofs for these statements can be found in Section B.1 and Section C.1. The accuracy and space analysis for algorithm 1 can be found in Appendix D.
Theorem 4.
  1. 1.
    (Privacy) BicriteriaDPCenters is $\varepsilon$-DP.
  2. 2.
    (Accuracy) With probability at least $1-k^{-\operatorname{poly}(k,\log\Lambda)}$,
  3. 3.
    (Space) BicriteriaDPCenters uses $\tilde{O}(k\operatorname*{poly}\left(\log\left({T}{\Lambda}k\right)\right))$ space.
  4. 4.
    (Size) $\mathcal{F}$} has at most $\tilde{O}({k}\log T)$ centers.
Recall that our main algorithm runs multiple concurrent instances of DP-Merge-Reduce at any given timestep, so $N$ in the theorem statement below represents the number of points seen by a specific instance of DP-Merge-Reduce. We use a non-DP coreset algorithm in DP-Merge-Reduce which we fix in the analysis to be the current state-of-the-art clustering result by[20] (see formal guarantees in Theorem 12).
Theorem 5 (DP-Merge-Reduce).
Let $0<\xi<1$, $T$ be the length of the entire stream, $\varepsilon,\delta$ be privacy parameters, $M$ be an arbitrary parameter such that $M>\frac{12}{\varepsilon}\log(\frac{2T}{\xi})$, and $P$ be a sub-stream of non-empty points with length $N$.
Suppose we are given black-box access to an offline $(\varepsilon_{1},\delta_{1})$-DP algorithm $\mathcal{A}$} that computes a $(\kappa,\eta_{1},\eta_{2})$-semicoreset of $X\subseteq{\mathbb{R}}^{d}$ of size $SZ_{\mathcal{A}}(N,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2},\xi_{A})$ using space $S_{\mathcal{A}}(N,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2},\xi_{A})$ with failure probability $\xi_{A}$. And we are given black-box access to an offline non-DP algorithm $\mathcal{B}$} that computes a $(1+\gamma)$-coreset of $X\subseteq{\mathbb{R}}^{d}$ of size $SZ_{{\mathcal{B}}}(N,k,d,\gamma,\xi_{B})$ using space $S_{\mathcal{B}}(N,k,d,\gamma,\xi_{B})$ with failure probability $\xi_{B}$. Then there exists an algorithm DP-Merge-Reduce (see algorithm 4) in the streaming model such that
  • (Privacy) DP-Merge-Reduce is $(\varepsilon+\varepsilon_{1},\delta_{1})$-DP.
[…]
  • (Space) DP-Merge-Reduce requires
    $S_{\mathcal{A}}(M,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2},\xi_{A})+S_{\mathcal{B}}(SZ_{\mathcal{A}}(M,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2},\xi_{A}),k,d,\gamma,\xi_{B})+\lceil\log(2N/M)\rceil\cdot S_{\mathcal{B}}(SZ_{\mathcal{B}}(M,k,d,\gamma,\xi_{B}),k,d,\gamma,\xi_{B})+3M/2$ space.
Remark.
The parameter $M$ denotes the block size of the base level in DP-Merge-Reduce (see algorithm 4). We treat $M$ as an arbitrary parameter in Theorem 5, Lemma 19 and Lemma 21. However we set $M$ to be a function of $\alpha$ (the multiplicative approximation error of the bicriteria solution), $\eta_{2}$ (the additive approximation error of the DP semicoreset algorithm) and a parameter $C_{M}$ in the proof of Theorem 13, i.e., $M:=\frac{\alpha\eta_{2}}{C_{M}}$. To obtain an $O(1)$-multiplicative approximation in Theorem 2, we assign $C_{M}=O(1)$. To obtain a $(1+\gamma^{\prime})$-multiplicative approximation in Theorem 3, we assign $C_{M}=\gamma/2$. See proof of Theorem 6 and Theorem 7 for details.
Privacy.
Lemma 1.
Proof.
[…]
2.1 Proofs of Theorem 1, Theorem 2 and Theorem 3
We now present the proof of the guarantees of our DP clustering framework (algorithm 1) stated in Theorem 1 below.
Proof.
The privacy guarantee follows from Lemma 1, and the accuracy guarantee follows from Theorem 14. The space guarantee follows from Lemma 19, where we simplify the second term $S_{\mathcal{B}}(SZ_{\mathcal{A}}(\cdot))$ in the statement of Lemma 19 by noting that the space used by the non-DP coreset algorithm $\mathcal{B}$} is linear in the semicoreset size of the algorithm $\mathcal{A}$}. ∎
Next we present the proofs of Theorem 2 and Theorem 3. We first restate the existing offline DP clustering and coreset results that we use as a blackbox from [47] and [32] in Theorem 6 and Theorem 7. Note that in Theorem 2, we need to convert the set of $k$ centers obtained from applying the DP clustering algorithm in Theorem 6 into a semicoreset via Theorem 11.
Theorem 6 (Theorem 3.10 in [47]).
The proof of Theorem 2 is stated below.
Proof.
We first note that using Theorem 11 together with Theorem 6 gives us an algorithm $\mathcal{A}$} that produces a $(\kappa,\eta_{1},\eta_{2})$-semicoreset where $\kappa=O(1),\eta_{1}=O(1)$ and $\eta_{2}={O}((k^{1.01}\cdot d^{0.51}+k^{1.5}+k)\cdot\operatorname*{poly}(\log(T)\log(\frac{1}{\delta}),\frac{1}{\varepsilon}))$ and that we can use as a black-box in our DP clustering framework. The privacy guarantee follows by observing that the resulting algorithm from Theorem 11 is $(3\varepsilon,\delta)$-DP and applying Item 1 of Theorem 1. We choose $C_{M}=100$ and the accuracy guarantee follows from Item 2 of Theorem 1.
Finally, for the space usage, we first determine the value of $M$ in Theorem 1. Because $\eta_{2}={O}((k^{1.01}\cdot d^{0.51}+k^{1.5})\cdot\operatorname*{poly}(\log(T)\log(\frac{1}{\delta}),\frac{1}{\varepsilon}))$ and $C_{M}=100$, we have that $M={O}((k^{1.01}\cdot d^{3.51}+d^{3}\cdot k^{1.5})\cdot\operatorname*{poly}(\log(T),\log(\frac{1}{\delta}),\frac{1}{\varepsilon}))$.
Note that $S_{\mathcal{A}}(M,k,d)=O((M+k)d)+O(k^{1+\eta}\cdot d\cdot\log(M))$ for some arbitrarily chosen constant $\eta>0$ where the second term comes from the space usage of Theorem 11. Also $SZ_{\mathcal{A}}(\cdot)=k$. The final space complexity is obtained by plugging in the value of $M$ and setting $\eta=0.01$. ∎
Theorem 7 (Lemma 16 in [32]).
There is a $2^{O_{\gamma}(d)}\operatorname*{poly}(n)$-time $\varepsilon$-DP algorithm that with probability 0.99 produces an $\left(1+\gamma,O_{\gamma}\left(\frac{k^{2}2^{O_{\gamma}(d)}}{\varepsilon}\operatorname*{poly}\log n\right)\right)$-coreset for $k$-means (and $k$-median). The size of the coreset is $2^{O_{\gamma}(d)}\cdot\operatorname*{poly}(k,\log n)$.
The proof of Theorem 3 is stated below.
Proof.
First note that the $\varepsilon$-DP coreset from Theorem 7 [32] is a $(\kappa,0,\eta)$-semicoreset where $\kappa=1+\gamma$ and $\eta=O_{\gamma}\left(\frac{k^{2}2^{O_{\gamma}(d)}}{\varepsilon}\operatorname*{poly}\log(T)\right)$ and thus we can use it directly as a blackbox in our DP clustering framework. The privacy guarantee follows from Item 1 of Theorem 1. The accuracy guarantee follows from Item 2 of Theorem 1 where we choose $C_{M}=\gamma/2$.
Finally, for the space usage, we first note that in Theorem 1, by plugging in the values of $C_{M}$ and $\eta$, we have that $M=O_{\gamma}\left(\frac{d^{3}k^{2}2^{O_{\gamma}(d)}}{\varepsilon}\operatorname*{poly}\log T\right)$. By observing that $S_{\mathcal{A}}(M,d,k)=\operatorname*{poly}(M,d,k)$ and total coreset size $SZ_{\mathcal{A}}(\cdot)=2^{O_{\gamma}(d)}\cdot\operatorname*{poly}(k,\log T)$ and plugging in the value of $M$ in Item 3 of Theorem 1, the space guarantee follows. ∎
3 Conclusion and Future Work
To the best of our knowledge, in this paper we designed the first sublinear space differentially private algorithms for $k$-clustering in the continual release setting (sublinear in stream length $T$). Our paper gives a general DP clustering framework that uses any existing offline DP clustering or coreset algorithm as a black-box and the space guarantee of the final algorithm depends on the additive error of the offline DP algorithm. In particular, we achieve a $(1+\gamma)$-multiplicative approximation using the DP coreset algorithm from [32] with $\tilde{O}_{\gamma}(\operatorname*{poly}(k,2^{O_{\gamma}(d)},\log(T)))$ space and an $O(1)$-multiplicative approximation with $\tilde{O}(k^{1.5}\cdot\operatorname*{poly}(d,\log(T)))$ space using the DP $k$-means algorithm from [47].
While our work achieves sublinear space in the insertion-only streaming setting, recently [39] studied the DP clustering problem under insertions and deletions but their results do not optimize for space — an exciting open direction is to design sublinear space differentially private algorithms for $k$-clustering in this model. Finally proving information-theoretic space lower bounds for any DP problem in the streaming setting is another exciting open direction. The only known space lower bound [22] is under cryptographic assumptions, and does not apply to clustering problems.
Appendix A Preliminaries
Norms and heavy hitters.